# Leveraging Soft Prompts for Privacy Attacks in Federated Prompt Tuning

### v1

@@ Abstract
Membership inference attack (MIA) poses a significant privacy threat in federated learning (FL) as it allows adversaries to determine whether a client’s private dataset contains a specific data sample. While defenses against membership inference attacks in standard FL have been well studied, the recent shift toward federated fine-tuning has introduced new, largely unexplored attack surfaces. To highlight this vulnerability in the emerging FL paradigm, we demonstrate that federated prompt-tuning, which adapts pre-trained models with small input prefixes to improve efficiency, also exposes a new vector for privacy attacks. We propose PromptMIA, a membership inference attack tailored to federated prompt-tuning, in which a malicious server can insert adversarially crafted prompts and monitors their updates during collaborative training to accurately determine whether a target data point is in a client’s private dataset. We formalize this threat as a security game and empirically show that PromptMIA consistently attains high advantage in this game across diverse benchmark datasets.  Our theoretical analysis further establishes a lower bound on the attack’s advantage which explains and supports the consistently high advantage observed in our empirical results. We also investigate the effectiveness of standard membership inference defenses originally developed for gradient or output based attacks and analyze their interaction with the distinct threat landscape posed by PromptMIA. The results highlight non-trivial challenges for current defenses and offer insights into their limitations, underscoring the need for defense strategies that are specifically tailored to prompt-tuning in federated settings.
@@ 1 Introduction
While Federated Learning (FL) avoids sharing raw data, it remains vulnerable to adversarial attacks (Muñoz-González et al., 2017; Zhu et al., 2019). One prominent class of attacks is membership inference attack (MIA)  (Shokri et al., 2017), where the adversarial server attempts to determine whether a specific data record was included in a client’s training dataset. Most existing studies largely focus on update based attacks, which leverage information such as gradients or model parameters exchanged during training to build an attack model  (Melis et al., 2019; Li et al., 2023). Recently, federated training of foundation models has shifted from full-model tuning to parameter-efficient fine-tuning (Han et al., 2024), where clients update only small subsets of parameters such as prompts (Wang et al., 2022) or adapters (Cai et al., 2022). In federated prompt tuning, clients optimize only small soft prompt vectors prepended to input embeddings. However, the effectiveness of MIAs in this new setting remains unclear, as the shared information is no longer the full model or gradients, but rather soft prompt vectors.
This shift in the form of shared updates introduces new challenges for existing FL based adversaries, which may no longer be applicable or effective against federated prompt tuning. At the same time, it opens up novel attack surfaces and raises important questions about the extent to which client-specific prompts can be exploited by an adversarial server to conduct privacy attacks. Previous work has examined privacy risks in LLM prompting, particularly in in-context learning (ICL), where discrete prompts (e.g., exemplars or task instructions) may leak sensitive information and have been shown vulnerable to eerence (Wen et al., 2024; Duan et al., 2023). For instance, adversaries can craft malicious queries such as targeted questions or carefully designed input–output pairs to test whether a particular record (e.g., a patient’s medical note) was included in the exemplars. To mitigate such risks, several defenses have been proposed for discrete prompts in ICL  (Wu et al., 2023; Hong et al., 2023; Tang et al., 2023). In contrast, the privacy risks of soft prompts, especially in federated prompt-tuning, remain largely unexplored. More importantly, the adversarial goal differs: in ICL based MIAs, the adversary aims to determine whether a specific exemplar was included in the discrete prompt context shown to the LLM. In federated prompt-tuning, the adversary instead exploits the clients’ learned prompt embeddings to infer whether a given training point was part of their local dataset.
To raise awareness of vulnerabilities in federated prompt tuning, we propose PromptMIA, a novel MIA specifically designed for the federated prompt tuning setting. PromptMIA leverages the prompt selection and update mechanism inherent to prompt learning frameworks. By injecting carefully crafted adversarial prompts into the shared prompt pool, the server can infer the presence of that sample based on whether the manipulated prompts are updated by the client. This attack achieves consistently high success rate across various benchmarks, highlighting a previously overlooked privacy vulnerability in prompt-based FL. Our contributions are as follows:
  • We propose PromptMIA, the first MIA specifically targeting federated prompt tuning. To the best of our knowledge, this is the first MIA to exploit soft prompts as an attack surface if FL (Sec. 3).
  • Unlike prior approaches that rely on gradients, model weights, or auxiliary datasets to train shadow models, PromptMIA exploits the prompt selection and update mechanism to infer membership in a single communication round, thereby reducing both computational cost and adversarial assumptions.
@@
  • We conduct extensive experiments across seven datasets (CIFAR-10, CIFAR-100, TinyImageNet, MNIST-M, Fashion-MNIST, CINIC-10, MMAFEDB) and three vision transformer architectures (ViT-B/32 (Dosovitskiy et al., 2020), DeiT-B/16 (Touvron et al., 2021), and ConViT (d’Ascoli et al., 2021)). Experimental results show that PromptMIA achieves over $90\%$ MIA success rate across diverse settings. (Sec. 4.1).
  • We provide a theoretical analysis that establishes a lower bound on the advantage of PromptMIA, which explains and supports its strong empirical performance across diverse settings (Sec. 3.2.3).
@@
  • Since PromptMIA exploits the prompt selection mechanism by monitoring which prompts are selected and updated rather than the content of the update, it naturally bypasses gradient obfuscation techniques such as DPSGD-based methods (Duan et al., 2023). Moreover, we show that even when the defender is aware of of the attack, it is not trivial to mitigate PromptMIA using existing techniques such as outlier detection or input noise perturbation (Sections 4.2 and 4.3). Together, these results expose the limitations of current countermeasures and underscore the urgent need for research on privacy defenses specifically tailored to federated prompt-tuning.
@@ 2 Federated Prompt-Tuning
In this section, we explain prompt-based learning and its extension to federated prompt tuning. Prompt-based learning (Lester et al., 2021) reformulates the downstream task adaptation problem as input modification rather than weight updates. Given an image $x\in\mathbb{R}^{H\times W\times C}$ and a pretrained ViT model with frozen embedding layer $f_{e}$, let $x_{e}=f_{e}(x)\in\mathbb{R}^{L_{x}\times D}$. $L_{x}$ is the number of patches and , and $D$ is the dimension of the patch embeddings. Prompt tuning prepends a learnable prompt $P_{e}\in\mathbb{R}^{L_{p}\times D}$ to $x_{e}$ to form $x_{p}=[P_{e};x_{e}]$. A frozen attention stack $f_{a}$ followed by classification head $f_{c}$ produces predictions $\hat{y}=f_{c}\!\big(f_{a}(x_{p})\big).$ Learning to Prompt (L2P) (Wang et al., 2022) maintains a prompt pool of size $M$, denoted as $\mathcal{P}=\{P_{i}\in\mathbb{R}^{L_{p}\times D}\}_{i=1}^{M}$ with corresponding learnable keys $\mathcal{K}=\{k_{i}\in\mathbb{R}^{D_{k}}\}_{i=1}^{M}$, $D_{k}$ is the dimension of each key vector. To facilitate query-key matching, a deterministic query function $q:\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{D_{k}}$ is used to encode input $x$ to the same dimension as the key. We use the pretrained model as a feature extractor and define the query feature as the [CLS] representation: $q(x)=f(x)[0,:].$ Using the cosine distance function $\gamma$, or inversely the cosine similarity function $\kappa$, the top-$N$ prompts are chosen using Eq. 1:
$\hat{\mathcal{K}}_{x}=\operatorname*{argmin}_{\{s_{i}\}_{i=1}^{N}\subseteq[M]}\sum_{i=1}^{N}\gamma\!\big(q(x),k_{s_{i}}\big)=\operatorname*{argmax}_{\{s_{i}\}_{i=1}^{N}\subseteq[M]}\sum_{i=1}^{N}\kappa\!\big(q(x),k_{s_{i}}\big),\qquad[M]=\{1,\dots,M\}.$ (1)
i.e., the $N$ prompts whose associated keys are closest to the query feature $q(x)$ under the cosine distance function $\gamma$. This is equivalent to choosing the $N$ keys with the highest cosine similarity to $q(x)$. The adapted input is $x_{p}=[P_{s_{1}};\cdots;P_{s_{N}};x_{e}].$ Let the average of the prompt-token hidden states be $\bar{h}(x_{p})=\mathrm{AvgPool}\!\big(f_{a}(x_{p})[0:NL_{p},:]\big)$ and the trainable classifier ${f_{c}}^{\phi}$ be parameterized by $\phi$, the final training objective is defined as:
@@
The first term is the softmax cross-entropy, while the second is a surrogate that pulls the selected keys to align with their corresponding query features. In federated prompt tuning, a server keeps a global pool $\mathcal{P}_{\textsc{global}}=\{P_{i}\}_{i=1}^{M}$. In each training round, given an input image $x$, the client selects $\hat{\mathcal{K}}_{x}$ using Eq. 1 and updates $\hat{\mathcal{K}}_{x}$, the corresponding $\hat{\mathcal{P}}_{x}$ along with ${f_{c}}^{\phi}$ according to Eq. 2. The server aggregates the trained prompts from participating clients to update global key-prompt pool $(\mathcal{K}_{\textsc{global}},\mathcal{P}_{\textsc{global}})$ using any of the existing FL algorithms (e.g., FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020)) or PFPT (Weng et al., 2024). See Appendix A.1 for more details.
@@ 3 MIAs against Federated Prompt-Tuning
In this section, we present the formulation and workflow of PromptMIA. An overview of the attack is shown in Fig. 1. Rather than analyzing model gradients or outputs like previous MIA approaches, PromptMIA operates by injecting adversarial prompts $\mathcal{P_{\textsc{adv}}}$ into the shared prompt pool $\mathcal{P_{\textsc{global}}}$ prior to a training round. These prompts are associated with adversarial key vectors $\mathcal{K_{\textsc{adv}}}$ tied to a target sample $\mathcal{T}$ and are designed to be activated only when $\mathcal{T}$ is present in a client’s data. As clients select and update prompts using query-key matching mechanism (Eq. 1) based on their local data, the server can monitor which injected prompts are modified and use this as a membership signal. This method leverages the prompt update mechanism as a covert privacy channel, enabling accurate membership inference.
Figure 1: PromptMIA workflow: (1) the server injects adversarial prompts designed for a target sample into the global pool; (2) the modified pool is broadcast to clients; (3) each client performs query-key matching, which selects all adversarial prompts if the target sample is present; (4) selected prompts are locally updated and (5) returned to the server. By monitoring which prompts are updated, the server infers the target’s membership in client data.
3.1 Prompt based AMI Threat Models as Security Games
@@
where $b$ is the true membership label (1 = in dataset, 0 = not), and $b^{\prime}$ is the adversary’s prediction. We further define the attack success rate (ASR) as:
$\textbf{ASR}^{\textup{AMI}}(\mathcal{A})=\frac{1}{2}(\Pr[b^{\prime}=1\mid b=1]+\Pr[b^{\prime}=0\mid b=0])=\tfrac{1}{2}\bigl(1+{\mathbf{\bf Adv}}^{\textup{AMI}}(\mathcal{A})\bigr).$ (4)
3.2 Membership Inference using Adversarial Prompt Injection
Given the security game laid out in Section 3.1, the adversary’s objective is to inject a set of $N$ adversarial keys $\mathcal{K}_{\textsc{adv}}$ together with their corresponding prompts $\mathcal{P}_{\textsc{adv}}$ into the global prompt pool such that, if the target data point $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected prompts will always be the adversarial set. The membership signal is defined as the event that all adversarial prompts are selected and subsequently updated. To build intuition, we first introduce a Naive Prompt Injection attack and analyze its weaknesses, before introducing the more robust method, PromptMIA.
3.2.1 Naive Prompt Injection Attack
In this naive approach, the adversary constructs adversarial keys $\mathcal{K}_{\textsc{adv}}=\{\,k_{a_{m}}\,\}_{m=1}^{N}$ and inserts them into the global prompt pool. The goal is to ensure that if target data $\mathcal{T}$ belongs to client’s dataset $\mathcal{D}$, the top-$N$ selected prompts using Eq. 1 always coincide with the adversarial set. A simple strategy is to align each adversarial key $k_{a_{m}}$ directly with the client’s query vector $q(\mathcal{T})$ so as to maximize cosine similarity. Let $\kappa$ be the cosine similarity operator. We ensure that:
$\kappa\!\left(q(\mathcal{T}),k_{a_{m}}\right)=(q(\mathcal{T})^{\top}k_{a_{m}})\;/\;\bigl(\|q(\mathcal{T})\|\cdot\,\|k_{a_{m}}\|\bigr)=1\;\text{ or }\;k_{a_{m}}=q(\mathcal{T}),\quad\forall\,k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}.$ (5)
One design choice is whether to inject new key–prompt pairs or modify existing ones. Injecting new pairs has two drawbacks: (1) clients could easily detect the sudden increase in the prompt pool size, and (2) there is no direct way to construct a matching prompt for a newly injected key. A more covert strategy is to modify existing key entries. The server selects a subset $S\subseteq[M]$ with $|S|=N$, and for each $j\in S$ performs key modificaton:
$(k_{j},P_{j})\;\mapsto\;(k_{a_{m}},P_{j}),\qquad m=1,\dots,N,$ (6)
However, this naive attack suffers from fundamental weaknesses. Since all adversarial keys collapse to the same vector $q(\mathcal{T})$ due to Eq. 5, clients can easily detect this redundancy through dimensionality reduction techniques (e.g., t-SNE, PCA) (see Fig. 2(b)), which would reveal a tight cluster of identical keys, or simply by directly inspecting key values. In addition, simple defenses such as discarding any key whose similarity exceeds a suspicious threshold (e.g., excluding $\kappa(q(\mathcal{T}),k_{a_{m}})>1-r$ for small $r>0$) can render the attack ineffective. Finally, because all adversarial keys are identical, the attack suffers from a high false positive rate: even when the target sample is absent, adversarial prompts may still be selected whenever the similarity of one adversarial key ( and as a result, all adversarial keys) to some non-target query $q(x)$, $x\in\mathcal{D},\,x\neq\mathcal{T}$, is higher than all benign keys, or $\kappa\!\left(q(x),k_{a_{m}}\right)>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(x),k_{b}\right)$.

### v2

@@ Abstract
Membership inference attacks (MIAs) pose a serious privacy threat in federated learning (FL). While MIAs have been extensively studied in standard FL, the recent shift toward federated fine-tuning introduces new and largely unexplored attack surfaces. In this work, we show that federated prompt-tuning, which adapts pre-trained models using lightweight input prefixes, exposes a novel and effective vector for membership inference. We propose PromptMIA, a membership inference attack tailored to federated prompt-tuning, in which a malicious server introduces adversarially crafted prompts and exploits their updates during collaborative training to determine whether a target data point belongs to a client’s private dataset. We formalize this threat via a security game and demonstrate that PromptMIA achieves consistently high attack advantage across diverse benchmark datasets. We also provide a theoretical lower bound on the attack advantage that explains the observed empirical behavior. Finally, we show that existing MIA defenses are often ineffective against PromptMIA, highlighting the need for defense mechanisms specifically tailored to prompt-tuning in federated settings.
Machine Learning, ICML
@@ 1 Introduction
Federated Learning (FL) (McMahan et al., 2017) is a prominent model training paradigm that enables data holders to collaboratively train a shared model without exposing their private data. However, while FL methods are privacy-compliant, they remain vulnerable to adversarial threats (Muñoz-González et al., 2017; Zhu et al., 2019). One prominent threat is membership inference attack (MIA)  (Shokri et al., 2017), where the adversarial server attempts to determine whether a specific data record was included in a client’s training dataset. Existing research has predominantly studied such privacy threats in settings where the attack surface is tied to clients’ gradients or full model parameters exchanged during training (Melis et al., 2019; Li et al., 2023). This classical view has, however, overlooked new attack surfaces that arise in more recent FL paradigms that instead communicate and aggregate external adaptation modules of pre-trained foundation models, e.g. soft-prompts prepended to input embeddings. (Weng et al., 2024).
To raise awareness of the potential existence of unexplored privacy threats in recent FL paradigms, this paper reveals a novel attack surface via manipulating the clustering and aggregation of local sets of soft prompts in federated prompt-tuning (FPT) (Weng et al., 2024). To the best of our knowledge, this is the first work that identifies an unexplored privacy risk in soft prompt optimization for representation fine-tuning, particularly in FPT. This is orthogonal to and complement existing research on privacy risks in prompting large language models with exemplars and/or task instructions that contain private data for in-context learning (ICL).
Such studies focus largely on designing adversarial queries that expose whether a particular data point was included in private instruction prompts used for in-context learning (Wen et al., 2024; Duan et al., 2023). In contrast, soft prompt optimization does not include private data in the prompts themselves. Moreover, the attack model in in-context learning originates from end users external to the training system whereas in federated prompt-tuning, it arises from the server within the training network. Existing defenses against prompt risk in in-context learning (Wu et al., 2023; Hong et al., 2023; Tang et al., 2023) are therefore not applicable to soft prompt risk in federated prompt-tuning.
To highlight the privacy risk that arises within soft prompt aggregation in FPT, we develop PromptMIA, a novel membership inference attack (MIA) that exploits FPT’s prompt update and selection mechanism as a new attack surface. This is substantiated with the following contributions:
(I) We develop an algorithm that optimizes adversarial prompts to be preferentially selected and updated by local clients when their datasets contain a target data point, while remaining unselected otherwise. PromptMIA is substantiated via running this algorithm on the server to create and add adversarial prompts to the shared prompt pools. The server can keep track of changes made to these prompts between communication iterations to infer membership information in a single communication round (Section 3.2.2).
(II) We analyze the theoretical attack advantage of PromptMIA from the lenses of a security game. In this view, the advantage of the attacker is characterized in terms of the true and false positive rates which measure the chance that the adversarial prompts are correctly and incorrectly selected, respectively. Our analysis shows that  PromptMIA achieves perfect true positive rate and provably low false positive rate, resulting in a high attack advantage (Section 3.2.3).
@@
(III) We run extensive experiments to validate the effectiveness of PromptMIA across $7$ datasets including CIFAR-10, CIFAR-100, TinyImageNet, MNIST-M, Fashion-MNIST, CINIC-10, MMAFEDB. We also run validation experiments across $3$ vision transformer architectures including ViT-B/32 (Dosovitskiy et al., 2020), DeiT-B/16 (Touvron et al., 2021), and ConViT (d’Ascoli et al., 2021). Our empirical results consistently show that PromptMIA achieves $90\%$ attack success rate, which corroborates the aforementioned high attack advantage (Section 4.1).
@@
(IV) We conduct additional analyses and experiments to demonstrate the effectiveness of PromptMIA against FL clients implementing standard defenses such as outlier detection, input noise perturbation, gradient obfuscation (Sections 4.2 and 4.3) and defenses that modify the prompt selection or aggregation protocol ( Appendix  A.4). We also note that PromptMIA exploits the prompt selection mechanism via monitoring which prompts are selected and updated rather than the content of the update, thus bypassing gradient obfuscation defenses (Duan et al., 2023). These results underscore the urgent need for more research on privacy defenses against this new prompt-based attack surface.
For clarity, we review the FPT framework in Section 2.
@@ 2 Federated Prompt-Tuning
In this section, we explain prompt-based learning and its extension to FPT. Prompt-based learning (Lester et al., 2021) reformulates the downstream task adaptation problem as input modification rather than weight updates. Given an image $x\in\mathbb{R}^{H\times W\times C}$ and a pretrained ViT model with frozen embedding layer $f_{e}$, let $x_{e}=f_{e}(x)\in\mathbb{R}^{L_{x}\times D}$. $L_{x}$ is the number of patches, and $D$ is the patch embedding’s dimension. A learnable prompt $P_{e}\in\mathbb{R}^{L_{p}\times D}$ is prepended to $x_{e}$ to form $x_{p}=[P_{e};x_{e}]$. A frozen attention stack $f_{a}$ followed by classification head $f_{c}$ produces predictions $\hat{y}=f_{c}\!\big(f_{a}(x_{p})\big).$ Learning to Prompt (L2P) (Wang et al., 2022) maintains a prompt pool of size $M$, denoted as $\mathcal{P}=\{P_{i}\in\mathbb{R}^{L_{p}\times D}\}_{i=1}^{M}$ with corresponding learnable keys $\mathcal{K}=\{k_{i}\in\mathbb{R}^{D_{k}}\}_{i=1}^{M}$. To facilitate query-key matching, a deterministic query function $q:\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{D_{k}}$ is used. We use the pretrained model as a feature extractor and define the query feature as the [CLS] representation: $q(x)=f(x)[0,:].$ Using the cosine distance function $\gamma$, the top-$N$ prompts are chosen using Eq. 1:
$\hat{\mathcal{K}}_{x}$ $=\operatorname*{argmin}_{\{s_{i}\}_{i=1}^{N}\subseteq[M]}\sum_{i=1}^{N}\gamma\!\big(q(x),k_{s_{i}}\big)$ (1)
i.e., the $N$ prompts whose associated keys are closest to the query feature $q(x)$ under the cosine distance $\gamma$, or equivaliently, the $N$ prompts whose associated keys have the highest cosine similarity to $q(x)$. The adapted input is $x_{p}=[P_{s_{1}};\cdots;P_{s_{N}};x_{e}].$ Let the average of the prompt-token hidden states be $\bar{h}(x_{p})$ and the trainable classifier ${f_{c}}^{\phi}$ be parameterized by $\phi$, the training objective is defined as:
@@
The first term is the softmax cross-entropy, while the second is a surrogate that pulls the selected keys to align with the corresponding query features. The server keeps a global prompt pool $\mathcal{P}_{\textsc{global}}=\{P_{i}\}_{i=1}^{M}$. Given input image $x$, the client selects $\hat{\mathcal{K}}_{x}$ using Eq. 1 and updates $\hat{\mathcal{K}}_{x}$, the associated prompts $\hat{\mathcal{P}}_{x}$ along with ${f_{c}}^{\phi}$ according to Eq. 2. The server aggregates the trained prompts from participating clients to update global key-prompt pool $(\mathcal{K}_{\textsc{global}},\mathcal{P}_{\textsc{global}})$ using any of the existing FL algorithms (e.g., FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020)) or PFPT (Weng et al., 2024). See Appendix A.1 for more details.
Figure 1: PromptMIA workflow: (1) the server injects adversarial prompts designed for a target sample into the global prompt pool; (2) the modified pool is broadcast to clients; (3) each client performs query-key matching, which selects all adversarial prompts if the target sample is present; (4) selected prompts are locally updated and (5) returned to the server. By monitoring which prompts are updated, the server infers the target’s membership in client data.
@@ 3 MIAs against Federated Prompt-Tuning
In this section, we present the formulation and workflow of PromptMIA. An overview of the attack is shown in Fig. 1. Unlike prior approaches that rely on gradients, model weights, or auxiliary datasets to train shadow models, PromptMIA operates by injecting adversarial prompts $\mathcal{P_{\textsc{adv}}}$ into the shared prompt pool $\mathcal{P_{\textsc{global}}}$ prior to a training round. These prompts are associated with adversarial key vectors $\mathcal{K_{\textsc{adv}}}$ and are designed to be selected only when target sample $\mathcal{T}$ is present in a client’s data. As clients select and update prompts (Eq. 1), the server can monitor which injected prompts are modified and use this as a membership signal. This method exploits prompt updates as a privacy channel to infer membership in a single round, reducing both computational cost and adversarial assumptions.
3.1 Prompt based AMIs as Security Games
@@
where $b$ is the true membership label, and $b^{\prime}$ is the adversary’s prediction. We define the attack success rate as:
$\textbf{ASR}^{\textup{AMI}}(\mathcal{A})$ $=\tfrac{1}{2}\bigl(1+{\mathbf{\bf Adv}}^{\textup{AMI}}(\mathcal{A})\bigr)$ (4)
$=\tfrac{1}{2}\bigl(\Pr[b^{\prime}=1\mid b=1]+\Pr[b^{\prime}=0\mid b=0]\bigr)$
@@
3.2 MIA using Adversarial Prompt Injection
Given the security game laid out in Section 3.1, the adversary’s objective is to inject a set of $N$ adversarial keys-prompts $(\mathcal{K}_{\textsc{adv}},\mathcal{P}_{\textsc{adv}})$ into the global prompt pool such that, if $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected prompts will always be the adversarial set. The membership signal is defined as the event that all adversarial prompts are selected and subsequently updated. To build intuition, we first introduce a Naive Prompt Injection attack and analyze its weaknesses, before introducing the more robust method, PromptMIA.
3.2.1 Naive Prompt Injection Attack
In this approach, the adversary constructs adversarial keys $\mathcal{K}_{\textsc{adv}}=\{\,k_{a_{m}}\,\}_{m=1}^{N}$ and inserts them into the global prompt pool. The goal is to ensure that if target data $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected prompts using Eq. 1 always coincide with $\mathcal{K}_{\textsc{adv}}$. A straightforward strategy is to align each adversarial key $k_{a_{m}}$ directly with the client’s query vector $q(\mathcal{T})$ so as to maximize cosine similarity. We ensure that:
$\kappa\!\left(q(\mathcal{T}),k_{a_{m}}\right)=1\;\text{or}\;k_{a_{m}}=q(\mathcal{T}),\quad\forall\,k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}.$ (5)
where $\kappa$ is the cosine similarity operator. The server selects a subset $S\subseteq[M]$ with $|S|=N$, and for each $j\in S$ performs key modificaton:
$(k_{j},P_{j})\;\mapsto\;(k_{a_{m}},P_{j}),\qquad m=1,\dots,N,$ (6)
However, this naive attack suffers from fundamental weaknesses. Since all adversarial keys collapse to the same vector $q(\mathcal{T})$ due to Eq. 5, clients can easily detect this redundancy through dimensionality reduction techniques (e.g., t-SNE, PCA) (see Fig. 2(b)), which would reveal a tight cluster of identical keys, or simply by directly inspecting key values. In addition, simple defenses such as discarding any key whose similarity exceeds a suspicious threshold (e.g., excluding $\kappa(q(\mathcal{T}),k_{a_{m}})>1-r$ for small $r>0$) can render the attack ineffective. Finally, because all adversarial keys are identical, the attack suffers from a high false positive rate ( see Section 4.1) : even when the target sample is absent, adversarial prompts may still be selected whenever the similarity of one adversarial key ( and as a result, all adversarial keys) to some non-target query $q(x)$, $x\in\mathcal{D},\,x\neq\mathcal{T}$, is higher than all benign keys, or $\kappa\!\left(q(x),k_{a_{m}}\right)>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(x),k_{b}\right)$.